# Normalization of Puiseux hypersurfaces

### v1

@@ Abstract.
It is known that the normalization of a quasi-ordinary complex singularity is a Hirzebruch-Jung, see [perez2000singularites, popescu2004analytical, aroca2005normal]. We extend this result to Puiseux hypersurfaces. Moreover, we prove that Hirzebruch-Jung singularities are precisely normalizations of Puiseux hypersurfaces. Our result holds over an algebraically closed field whose characteristic does not divide the degree of the polynomial defining the hypersurface. Finally, in the analytic complex case, we conclude that the normalization of an irreducible Puiseux hypersurface is the normalization of a complex analytic quasi-ordinary singularity.
@@
This work has been partially supported by Laboratorio Internacional Solomon Lefschetz, by DGAPA- Universidad Nacional Autónoma de México through PAPIIT IN113323, and by SECIHTI, Mexico CVU’s: 1337652, 781791, 1044169.
We are grateful to M. Spivakovsky for his helpful comments and suggestions, which greatly improved this work. The corresponding author also thanks P. Popescu-Pampu and P. González-Pérez for providing valuable references.
@@ Introduction
An analytic germ of dimension $n$ is quasi-ordinary when it is a local covering of $\mathbb{K}^{n}$ unramified outside the coordinate hyperplanes. In the complex analytic case, the Abhyankar–Jung theorem [abhyankar1955ramification] states that quasi-ordinary hypersurfaces are Puiseux. Quasi-ordinary singularities have been widely studied; see, for example, [alonso1989algorithm, MR2957197, mourtada2015polyhedral]. For an introductory text, we refer to [da2011semigrupo]. Hironaka defined a more general class of singularities containing the quasi-ordinary ones, which he called $\nu$–quasi-ordinary [hironaka2003theory]; see also [cassou2011nu, aroca2024nu].
It is natural to ask which properties can be extended from quasi-ordinary to Puiseux hypersurfaces. For instance, Lipman and Gau [lipman1988topological] associated to a root of a quasi-ordinary polynomial an ordered set of $n$-tuples of nonnegative rational numbers, called characteristic exponents. Let $\lambda^{(1)},\ldots,\lambda^{(g)}$ be the set of characteristic exponents of a quasi-ordinary root $\xi$. Then
@@
In [tornero2008kummer], Tornero proved that if $\xi$ is a root of a Puiseux polynomial, there exist $\lambda^{(1)},\ldots,\lambda^{(g)}$ such that (1) holds. Such a set is called a set of distinguished exponents of $\xi$. In the quasi-ordinary case, it is known that the characteristic exponents determine the topological type of the singularity. The relationship between the topological type of a Puiseux hypersurface and its distinguished exponents is not yet known.
One of the most important properties of a quasi-ordinary singularity is that its normalization is Hirzebruch–Jung, that is, analytically isomorphic to the zero-dimensional orbit of an affine toric variety defined by a maximal simplicial pair [perez2000singularites, aroca2005normal, popescu2004analytical]. This paper is devoted to showing that this property also holds for Puiseux hypersurfaces (see Theorem 5.2).
@@
As a corollary, using a result of Popescu-Pampu [popescu2003higher], in the complex analytic case we conclude that the normalization of an irreducible Puiseux hypersurface coincides with the normalization of a complex analytic quasi-ordinary singularity (see Corollary 5.6).
@@ 1. Affine Semigroups
See [cox2024toric, Proposition 1.2.17] and [Kempf, page 7].
@@
See [cox2024toric, Proposition 1.3.8] and [Kempf, Theorem 1’, Chapter 1, Section 1].
@@ 2. Normal Toric Varieties
See [cox2024toric, Propositions 1.1.14 and 1.3.5], [Kempf, Theorem 1, Chapter 1, Section 1].
@@
See  [cox2024toric, Propositions 1.3.5 and 1.3.8],  [Kempf, Theorem 1, Chapter 1, Section 1].
@@ 3. Analytical Equivalence
The completion of $\mathbb{K}[S]$ with respect to its maximal ideal $\mathfrak{m}=\langle X^{\alpha}\mid\alpha\in S,\alpha\neq 0\rangle$ is $\mathbb{K}[[S]]$, see [eisenbud2013commutative, Example 10.14, p. 181]. Since excellent rings satisfy that normalization and completion commute, see [egaiv1964grothendieck, Proposition 7.8.3.(vii)], it follows from Lemma 2.4 that
@@
For example (Hartshorne, Example 5.6.3 [hartshorne2013algebraic]), let $\mathbb{X}=V(x^{2}(x+1)-y^{2})$ and $\mathbb{Y}=V(xy)$. Then the singularity $(0,0)$ in $\mathbb{X}$ is analytically isomorphic to $(0,0)$ in $\mathbb{Y}$. Geometrically, near the origin both are the union of two nonsingular branches with distinct tangents.
@@ 4. Puiseux Hypersurfaces
Let $\xi\in\mathbb{K}[[X^{*}]]$. $E=\left\{\lambda^{(1)},\ldots,\lambda^{(g)}\right\}\subset\text{Supp}\left(\xi\right)$ is a set of distinguished [^1] Given a Puiseux series $\xi\in\mathbb{K}[[X^{*}]]$, in [tornero2008kummer] the elements of a set $\left\{k\lambda^{(1)},\ldots,k\lambda^{(g)}\right\}\subset k\ \text{Supp}\left(\xi\right)\subset(\mathbb{Z}_{\geq 0})^{n}$ such that $\lambda^{(1)},\ldots,\lambda^{(g)}$ satisfy the equation (2) are called distinguished exponents. exponents of $\xi$ if
@@
In [tornero2008kummer], Tornero proves that all Puiseux series $\xi$, with minimal polynomial $f$ such that char $(\mathbb{K})$ does not divide  $\text{ deg }(f)$, has a set of distinguished exponents.
@@
for $i=2,\ldots,g$ with $g=\min\{i\ |\ \lambda^{(i)}\neq\infty\}.$ For more details we refer to [tornero2008kummer].
@@ 5. Main Results
Lipman proved in the complex case that every normal quasi-ordinary singularity is isomorphic to a germ of the form $y^{k}-x_{1}\cdots x_{n}$ for $k\geq 2$, see [lipman1988topological, Remark 7.3.2]. In other words, every normal quasi-ordinary singularity is analytically isomorphic to the affine toric variety $\mathrm{Spec}(\mathbb{C}[S])$, where $S$ is the affine semigroup generated by $(1/k,\ldots,1/k)$ and $(\mathbb{Z}_{\geq 0})^{n}$.
@@
Its characteristic exponents are $(3/2,1,0)$ and $(2,3/2,1)$, see [Ayala2025]. The saturation of the affine semigroup $S$ generated by these exponents and $(\mathbb{Z}_{\geq 0})^{3}$ is $\frac{1}{2}\mathbb{Z}_{\geq 0}\times\frac{1}{2}\mathbb{Z}_{\geq 0}\times\mathbb{Z}_{\geq 0}\simeq(\mathbb{Z}_{\geq 0})^{3}$. Therefore
@@
The roots of $f=(z^{2}-x-y)^{2}-4xy\in\mathbb{K}[[x,y]][z]$ are the Puiseux series $\pm x^{1/2}\pm y^{1/2}$, which are not quasi-ordinary, see [Ayala2025]. Following Tornero’s algorithm, we obtain distinguished exponents $(1/2,0)$ and $(0,1/2)$. The affine semigroup $S$ generated by $(\mathbb{Z}_{\geq 0})^{2}$ and these exponents is already saturated. The normalization of the hypersurface defined by $f$ is analytically isomorphic to the non-singular toric surface $\mathbb{K}[S]\simeq\mathbb{K}^{2}$.
In the complex-analytic case, Popescu-Pampu [popescu2003higher, Proposition 3.5] proved that Hirzebruch-Jung singularities are precisely the normal quasi-ordinary singularities. Hence, we have:
@@
Declarations
Competing Interests: The authors have no relevant financial or non-financial interests to disclose.
Affiliations

### v2

@@ Abstract.
It is known that the normalization of a quasi-ordinary complex singularity is a Hirzebruch-Jung, see [11, 18, 3]. We extend this result to Puiseux hypersurfaces. Moreover, we prove that Hirzebruch-Jung singularities are precisely normalizations of Puiseux hypersurfaces. Our result holds over an algebraically closed field whose characteristic does not divide the degree of the polynomial defining the hypersurface. Finally, in the analytic complex case, we conclude that the normalization of an irreducible Puiseux hypersurface is the normalization of a complex analytic quasi-ordinary singularity.
@@
2020 Mathematics Subject Classification: 32S05, 14B05, 13F25, 14M25.
  This work has been partially supported by Laboratorio Internacional Solomon Lefschetz, by DGAPA- Universidad Nacional Autónoma de México through PAPIIT IN113323, and by SECIHTI, Mexico CVU’s: 1337652, 781791, 1044169. We are grateful to M. Spivakovsky for his helpful comments and suggestions, which greatly improved this work. The corresponding author also thanks P. Popescu-Pampu and P. González-Pérez for providing valuable references.
@@ Introduction
An analytic germ of dimension $n$ is quasi-ordinary when it is a local covering of $\mathbb{K}^{n}$ unramified outside the coordinate hyperplanes. In the complex analytic case, the Abhyankar–Jung theorem [1] states that quasi-ordinary hypersurfaces are Puiseux. Quasi-ordinary singularities have been widely studied; see, for example, [2, 7, 17]. For an introductory text, we refer to [9]. Hironaka defined a more general class of singularities containing the quasi-ordinary ones, which he called $\nu$–quasi-ordinary [14]; see also [5, 4].
It is natural to ask which properties can be extended from quasi-ordinary to Puiseux hypersurfaces. For instance, Lipman and Gau [16] associated to a root of a quasi-ordinary polynomial an ordered set of $n$-tuples of nonnegative rational numbers, called characteristic exponents. Let $\lambda^{(1)},\ldots,\lambda^{(g)}$ be the set of characteristic exponents of a quasi-ordinary root $\xi$. Then
@@
In [20], Tornero proved that if $\xi$ is a root of a Puiseux polynomial, there exist $\lambda^{(1)},\ldots,\lambda^{(g)}$ such that (1) holds. Such a set is called a set of distinguished exponents of $\xi$. In the quasi-ordinary case, it is known that the characteristic exponents determine the topological type of the singularity. The relationship between the topological type of a Puiseux hypersurface and its distinguished exponents is not yet known.
One of the most important properties of a quasi-ordinary singularity is that its normalization is Hirzebruch–Jung, that is, analytically isomorphic to the zero-dimensional orbit of an affine toric variety defined by a maximal simplicial pair [11, 3, 18]. This paper is devoted to showing that this property also holds for Puiseux hypersurfaces (see Theorem 5.2).
@@
As a corollary, using a result of Popescu-Pampu [19], in the complex analytic case we conclude that the normalization of an irreducible Puiseux hypersurface coincides with the normalization of a complex analytic quasi-ordinary singularity (see Corollary 5.6).
@@ 1. Affine Semigroups
See [8, Proposition 1.2.17] and [15, page 7].
@@
See [8, Proposition 1.3.8] and [15, Theorem 1’, Chapter 1, Section 1].
@@ 2. Normal Toric Varieties
See [8, Propositions 1.1.14 and 1.3.5], [15, Theorem 1, Chapter 1, Section 1].
@@
See  [8, Propositions 1.3.5 and 1.3.8],  [15, Theorem 1, Chapter 1, Section 1].
@@ 3. Analytical Equivalence
The completion of $\mathbb{K}[S]$ with respect to its maximal ideal $\mathfrak{m}=\langle X^{\alpha}\mid\alpha\in S,\alpha\neq 0\rangle$ is $\mathbb{K}[[S]]$, see [10, Example 10.14, p. 181]. Since excellent rings satisfy that normalization and completion commute, see [12, Proposition 7.8.3.(vii)], it follows from Lemma 2.4 that
@@
For example (Hartshorne, Example 5.6.3 [13]), let $\mathbb{X}=V(x^{2}(x+1)-y^{2})$ and $\mathbb{Y}=V(xy)$. Then the singularity $(0,0)$ in $\mathbb{X}$ is analytically isomorphic to $(0,0)$ in $\mathbb{Y}$. Geometrically, near the origin both are the union of two nonsingular branches with distinct tangents.
@@ 4. Puiseux Hypersurfaces
Let $\xi\in\mathbb{K}[[X^{*}]]$. $E=\left\{\lambda^{(1)},\ldots,\lambda^{(g)}\right\}\subset\text{Supp}\left(\xi\right)$ is a set of distinguished [^1] Given a Puiseux series $\xi\in\mathbb{K}[[X^{*}]]$, in [20] the elements of a set $\left\{k\lambda^{(1)},\ldots,k\lambda^{(g)}\right\}\subset k\ \text{Supp}\left(\xi\right)\subset(\mathbb{Z}_{\geq 0})^{n}$ such that $\lambda^{(1)},\ldots,\lambda^{(g)}$ satisfy the equation (2) are called distinguished exponents. exponents of $\xi$ if
@@
In [20], Tornero proves that all Puiseux series $\xi$, with minimal polynomial $f$ such that char $(\mathbb{K})$ does not divide  $\text{ deg }(f)$, has a set of distinguished exponents.
@@
for $i=2,\ldots,g$ with $g=\min\{i\ |\ \lambda^{(i)}\neq\infty\}.$ For more details we refer to [20].
@@ 5. Main Results
Lipman proved in the complex case that every normal quasi-ordinary singularity is isomorphic to a germ of the form $y^{k}-x_{1}\cdots x_{n}$ for $k\geq 2$, see [16, Remark 7.3.2]. In other words, every normal quasi-ordinary singularity is analytically isomorphic to the affine toric variety $\mathrm{Spec}(\mathbb{C}[S])$, where $S$ is the affine semigroup generated by $(1/k,\ldots,1/k)$ and $(\mathbb{Z}_{\geq 0})^{n}$.
@@
Its characteristic exponents are $(3/2,1,0)$ and $(2,3/2,1)$, see [6]. The saturation of the affine semigroup $S$ generated by these exponents and $(\mathbb{Z}_{\geq 0})^{3}$ is $\frac{1}{2}\mathbb{Z}_{\geq 0}\times\frac{1}{2}\mathbb{Z}_{\geq 0}\times\mathbb{Z}_{\geq 0}\simeq(\mathbb{Z}_{\geq 0})^{3}$. Therefore
@@
The roots of $f=(z^{2}-x-y)^{2}-4xy\in\mathbb{K}[[x,y]][z]$ are the Puiseux series $\pm x^{1/2}\pm y^{1/2}$, which are not quasi-ordinary, see [6]. Following Tornero’s algorithm, we obtain distinguished exponents $(1/2,0)$ and $(0,1/2)$. The affine semigroup $S$ generated by $(\mathbb{Z}_{\geq 0})^{2}$ and these exponents is already saturated. The normalization of the hypersurface defined by $f$ is analytically isomorphic to the non-singular toric surface $\mathbb{K}[S]\simeq\mathbb{K}^{2}$.
In the complex-analytic case, Popescu-Pampu [19, Proposition 3.5] proved that Hirzebruch-Jung singularities are precisely the normal quasi-ordinary singularities. Hence, we have: